OpenStudy (anonymous):

Solve for x. 12 – 3x – x = 8 A. –10 B. –5 C. 1 D. 2

hartnn (hartnn):

-3x-x factoring out x x (-3-1) = ... ?

OpenStudy (anonymous):

-4

hartnn (hartnn):

-4x you mean? yes, so we have 12 -4x = 8 subtract 12 on both sides, what u get ?

OpenStudy (anonymous):

A

hartnn (hartnn):

its not A -4x = 8-12 -4x = -4 now can you find x ?

hartnn (hartnn):

just divide both sides by -4

hartnn (hartnn):

x = (-4)/(-4) = 1